if an oxide of carbon is 20% by mass what is its empirical formula




1
Expert's answer
2021-03-12T05:57:01-0500

Let's assume that the compound has a mass of 100g.


Moles of Carbon ="\frac{80g}{12g/mol}=6.66mol"


Moles of Oxygen ="\frac{20}{16g/mol}=1.25mol"


"C:\frac{6.66}{6.66}=1"


"O:\frac{1.25} {6.66}=0"


Thus the imperial formula "=CO"